calculate the mass of sodium oxide formed whem 3.45g of sodium burns in air
1
Expert's answer
2018-06-20T05:21:26-0400
4Na + O2 → 2Na2O

n(Na) = m/M = 3.45g/23g/mol = 0.15 mol

n(Na2O) = n(Na)/2 = 0.15/2 = 0.075 mol

M(Na2O) = 23*2 + 16 = 62 g/mol

m(Na2O) = n*M = 0.075mol*62g/mol = 4.65 g